OpenStudy (anonymous):

If a fox starts from rest and accelerates at a rate of 1.1 m/s2, how long does it take the fox to run 5 m?

OpenStudy (anonymous):

You need to use a linear formula We have u,v,a,t,s u = 0 a = 1.1 s = 5 we want to find t The formula we need is s=ut+0.5at^2 5 = 0 + 0.5*1.1*t^2 5 = 0.55 t^2 t = sqrt(5/0.55) t = 9.09 seconds Therefore it will take 9.09 seconds

OpenStudy (anonymous):

it took 3.0s

OpenStudy (anonymous):

Sorry, forgot to take the square root. 3 seconds it is. :P

OpenStudy (anonymous):

A runner starts from rest and has an acceleration of 3 m/s2. How fast is she running after 1.1 seconds?

OpenStudy (anonymous):

@Riddellikins

OpenStudy (anonymous):

3*1.1 = 3.3m/s

OpenStudy (anonymous):

A truck is moving 14 m/s to the south. If it takes the truck 20 seconds to reduce its velocity to 10 m/s south, what is its acceleration?

OpenStudy (anonymous):

4/20 = 0.2m/s/s

OpenStudy (anonymous):

north or south?

OpenStudy (anonymous):

north

OpenStudy (anonymous):

A car crash causes an acceleration of 6.2 g. What is the car's acceleration in meters per second squared?

OpenStudy (anonymous):

6.2*9.81=60.822

OpenStudy (anonymous):

A runner starts from rest and speeds up with constant acceleration. If she has gone a distance of 30 m at the point when she reaches a speed of 8 m/s, what is her acceleration?

OpenStudy (anonymous):

u = 0 v = 8 a = ? s = 30 What formula do you need?

OpenStudy (anonymous):

acceleraton solving for velocity

OpenStudy (anonymous):

But what is the actual formula?

OpenStudy (anonymous):

doesnt say choices are: 1.1 m/s 2 3.8 m/s 2 9.1 m/s 2 7.5 m/s 2

OpenStudy (anonymous):

Yeah, you have to know the formula yourself. There are four of them (for linear motion under a constant acceleration) v = u +at v^2 = u^2 + 2as s = ut + 0.5at^2 s = 0.5(u+v)t

OpenStudy (anonymous):

a= Change in velocity / change in time

OpenStudy (anonymous):

No, it's one of the ones that I typed out. Have you learnt them?

OpenStudy (anonymous):

no but it was 1.1

OpenStudy (anonymous):

A cheetah accelerates from rest at a rate of 5.4 m/s2. How far has the cheetah gone after 1.8 seconds